# Generating Synthetic Data via Augmentations for Improved Facial Resemblance in DreamBooth and InstantID

### v1

@@ Abstract
The personalization of Stable Diffusion for generating professional portraits from amateur photographs is a burgeoning area, with applications in various downstream contexts. This paper investigates the impact of augmentations on improving facial resemblance when using two prominent personalization techniques: DreamBooth and InstantID. Through a series of experiments with diverse subject datasets, we assessed the effectiveness of various augmentation strategies on the generated headshots’ fidelity to the original subject. We introduce FaceDistance, a wrapper around FaceNet, to rank the generations based on facial similarity, which aided in our assessment. Ultimately, this research provides insights into the role of augmentations in enhancing facial resemblance in SDXL-generated portraits, informing strategies for their effective deployment in downstream applications.
Figure 1: Pipeline for creating personalized images based on synthetically generated images through classical and GenAI-based augmentations for better downstream resemblance in DreamBooth-generated images.
@@ 1 Introduction
In this work, we analyze the effect of augmentations on personalized SD models trained with few-shot and zero-shot methods, particularly DreamBooth and InstantID. We investigate how different augmentations impact model performance and whether they enhance the realism and consistency of generated images.
In particular, we analyze both classical and generative augmentation strategies to bridge the gap between limited real data and high-fidelity synthetic outputs. By refining facial features and preserving identity through targeted GenAI-based augmentations, such as InstantID, we aim to improve the applicability of personalized generation in scenarios where synthetic data must closely mirror real-world characteristics. We analyze under which conditions we can ensure that ”GenAI outputs improve GenAI outputs”, avoiding a data quality collapse, providing best practices and heuristics.
@@ 2.2 Subject-Driven Image Generation
InstantID [20] is a zero-shot method that combines facial feature extraction with text conditioning. It extracts five key facial landmarks to condition the position and orientation of the generated face, providing greater control over the output.
We standardize our experiments using the same SDXL model for both techniques to ensure fair methodological comparison.
@@ 2.3.1 Classical Image Augmentations
Classical image augmentation techniques include geometric transformations (flipping, rotation, scaling, cropping), photometric adjustments (brightness, contrast, saturation, hue), and noise injections (Gaussian, salt-and-pepper). These predefined transformations maintain semantic integrity while introducing controlled diversity to expand limited training datasets.
@@ 2.4 Face Processing Approaches
FaceNet [17] maps facial images to a 128-dimensional embedding space where similar faces are positioned closely together. The standard pipeline uses MTCNN [22] for face detection before embedding generation, with cosine distance metrics for similarity assessment [18].
@@
Figure 2: Pipeline for creating Gen-AI augmented personalized data via InstantID. Based on one or more input images of a person, we run it through the InstantID pipeline but with augmented landmarks and prompts. The landmarks are taken from the input image and slightly perturbed for good resemblance. The collected synthetic dataset is then further used for downstream DreamBooth training. Figure modified from [20].
@@ 3 Methodology
3.1 Subject Datasets.
Our dataset consists of 3 to 15 images per participant, with $n=10$ participants. To maintain a naturalistic data collection process, we instructed them: “Can you send me portrait/selfie-style photos of your face in different places? The more different places, the better.” By avoiding rigid guidelines, we ensured that the collected images reflect realistic user behavior. As a result, our findings are well-aligned with real-world data distributions, enhancing the transferability and applicability of our results.
Our dataset exhibits a diverse range of environmental conditions, facial orientations, and image qualities, ensuring variability that mirrors real-world scenarios. The images encompass different backgrounds, lighting conditions, and subject behaviors, contributing to a dataset that is both representative and robust. For instance, some images feature cluttered or irregular backgrounds (e.g., Baker-Zoe, Bottle-Hugo), while others are captured in controlled environments (Biometric-Kora). Variation in gaze direction is also present, with Doctor-Nina not looking at the camera, while 3D-Gary includes dynamic head movements extracted from a video. Additionally, differences in personal appearance and accessories are observed, such as Farmer-Lisa wearing a helmet and Staircase-Judy wearing makeup. Lighting conditions range from well-lit (Vacation-Anna) to suboptimal (2024-Kora), further enhancing the dataset’s realism. These characteristics make our dataset a valuable resource for evaluating model performance under unconstrained, real-world conditions.
@@ 3.2 Dataset Augmentations
Classical Augmentations  Standard techniques include: (i) Random Horizontal Flip with $p\in\{0,\frac{1}{2},1\}$, and (ii) Color Jitter varying brightness, contrast, saturation (±$(\pm 5,\pm 15)$ and hue (±$(\pm 5\degree)$.
Background Augmentation  We use U2-Net [14] for subject isolation, testing both base and human segmentation models. Backgrounds include flat colors, patterns from Wikimedia [5], and Flickr Places.
@@
Generative Augmentation  Using InstantID, we generate new subject images with prompts from dolphin 2.2.1 - Mistral 7B [3] and varied facial landmarks (Figure 2).
3.3 Hardware, Software, and Hyperparamters
@@
Default prompt
we use “A professional headshot of a subject wearing a suit standing in a well-lit studio, DSLR” as the default prompt. Empirical evidence suggests including the gender as man and woman gives generated images gender characteristics based on western culture, which we preferred.
@@ 3.4 FaceDistance Metric
For a subject image dataset, we calculate their embedding using FaceNet [17], which maps similar faces to similar locations on a hypersphere. Then, we calculate the mean face vector $v_{real}$. For a given generated face, we measure its embedding vectors cosine distance to $v_{real}$.
FaceDistance is a useful technique for distinguishing between ”good” and ”bad” generations. This can be used to rank generated images based on their similarity (lower is better). It can be used to discard the largest nof distances to improve personalization pipelines.
For our subject datasets, the mean cosine distance of $v_{real}$ to real images is $\bar{v}_{within\ real}\approx 0.13$. We notice $\max(v_{within\ real})=0.22$ and $\min(v_{within\ real})=0.05$.
@@ 4 Experiment Results
Despite selecting a realistic image generation model, achieving photorealistic generation of an individual’s face remains challenging without imposing strict constraints on the subject images. We have relaxed many of these constraints to enhance usability, as expecting an average user to compile a dataset of themselves without understanding the underlying image generation techniques presents a significant challenge. Ensuring high subject fidelity is crucial for these methods to be effective in downstream applications, as humans are highly sensitive to variations in facial features compared to textures.
One major issue with datasets without great constraints is that the images is not a good representation of the person. It can be compared to having difficulty recognizing a person in real life whom you only saw in photographs. We observe this phenomenon for small datasets with $\operatorname{size}\leq 3$. In these cases, the generated images is a good reflection of the dataset (if someone doesn’t know them in real life, they are likely to claim that these pictures are good. Otherwise the generated images are not a good representation of the real person).
(a) Real Images
Figure 3: Example improvement of including Instantid generated images in the subject dataset Vacation-Anna. The model is prompted with default prompt with batchsize 4. The results are not cherry-picked to resemble the downstream application use. Although (b) is visually more interesting, the method in (c) is more consistent across many subject datasets.
@@ 4.1 DreamBooth
We configure our hyperparameters such that recontextualization capabilities can be sacrificed for high facial fidelity. Identity preservation is hard in DreamBooth. so we rather overfit to achieve high subject fidelity and have limited freedom in generations.
@@
The common theme in augmentations is that if the augmented image has any kind of artifact/anomaly, then the rare token will be associated with it. The supporting observations are (i) When background is replaced with a geometric pattern (from wikimedia patterns), the model will focus on learning the pattern than the subject (ii) When image is upscaled with ESRGAN, the texture ESRGAN introduces say present in generations (iii) the masks generated with U2-Net is not pixel-perfect. and results in a mix around hair/air boundary. This mix becomes associated with the subject. The human segmentation models training data was not highly accurate around hairs but was better in identifying body parts. The base model is performs better around hair and was overall better. The robustness of human segmentation model is not needed. (iii) any kind of color jitter is visible in generated images. For example the saturation change of $0.1$ is clearly present in generations. (iv) using Adobe Lightroom as a preprocessing step resulted in better color graded generations compared to non-preprocessed datasets. (v) datasets with low contrast (e.g. exclusively Polaroid pictures) resulted in copying the photography style/lighting from the pictures — though this can be attributed to our hyperparameter configuration.
@@
Because of the low recontextualization capabilities, backgrounds becomes highly associated with the rare token. Replacing the background with Pastel Colors and Rainbow Colors led to eccentric and often unrealistic images, with the latter occasionally generating pictures without subjects. Gray offered the highest resemblance to the subject, while Dark Gray caused the model to disassociate the subject from its context. Because of problems with U2-Net, Light Gray background outperformed Dark Gray, especially in bright environments.Wikimedia Patterns slowed down learning and degraded the image quality across all generations. Lastly, Studio Backdrops introduced irregularities that reduced the quality of the generated images which can be thought as similar to Wikimedia Patterns because backdrops has patterns.
@@
Random Horizontal Flip slowed learning due to face asymmetry, which confused facial features. Random Rotation caused distorted images and introduced black padding bars, which also can be seen in augmented subject images. Color Jitter led to undesirable results, as brightness, contrast, saturation, and hue changes were linked to rare tokens, causing erratic generations.
@@
Both Alpha Blending and Poison Blending are discouraged, as they require careful manual processing to achieve good results. These techniques are not straightforward to apply and can lead to undesirable artifacts if not handled properly.
@@
Images around 1 Megapixel performed best, providing a balanced resolution for high-quality generation. Upscaling with ESRGAN introduced visible artifacts, especially around facial features. Upscaling with Lanczos was effective, particularly when starting from larger images. However, if the initial dataset contained low-resolution images, the generated images exhibited facial blurring due to the nature of the Lanczos algorithm. The difference between bicubic and Lanczos was negligible. Downscaling resulted in lower-quality generations compared to using original-sized images. It should be noted that our testing output resolution was $1024\times 1024$.
InstantID Augmentation
@@
Datasets augmented with InstantID yield clearly superior performance. The added images need to be diverse (i.e., generated with various text conditioning and different keypoint images). Since we trade recontextualization abilities for increased facial similarity, generating the same person in similar contexts is beneficial. DreamBooth achieves similar facial similarity compared to InstantID but allows for greater control. The rigidity caused by the keypoint images is eliminated. However, this method is more computationally expensive than raw InstantID. Additionally, achieving proper prompt diversity can be challenging. I prefer InstantID over DreamBooth.
@@
The ratio of real to InstantID-generated images depends entirely on the diversity of the generated images. One rule of thumb is that no single concept should comprise more than 25% of the dataset. For example, if images labeled as ”a [V] man in a library” exceed 25%, DreamBooth training will associate the rare token with the concept. This results in a final DreamBooth model that is unusable due to a complete loss of recontextualization ability caused by overfitting.
@@
Since InstantID generations are highly realistic, one can generate additional images with it to better represent the subject during DreamBooth training. We use the same diffusion model for both InstantID and DreamBooth to integrate the subject more effectively into the model without altering the subject’s context. This ensures that the dataset distribution remains closer to the diffusion model’s generation space.
Figure 4: FaceDistance Distributions of 2000 Samples from Different Saved Dreambooth finetunes of SDXL Real. Closeup-Kora is used. The KDE for each looks like a normal distribution.
@@ 4.1.1 FaceDistance
We tried to select the “best” DreamBooth checkpoint by generating images of “a [V] man” in different contexts for all checkpoints and ranking them using FaceDistance. This method was able to discard obviously bad checkpoints (e.g. anomalies in generations, unable to generate the subject, divergence) but is not able to rank “good enough” checkpoints within themselves. (Figure 4 The FaceNet manifold isn’t sensitive to very similar looking people. For a given hyperparameter configuration, a few tests show when the model will be converged to its best state (usually between 3k and 6k steps) and since FaceDistance isn’t able to differentiate betweent them, FaceDistance isn’t a useful tool for this purpose.

### v2

@@ Abstract
Personalizing Stable Diffusion for professional portrait generation from amateur photos faces challenges in maintaining facial resemblance. This paper evaluates the impact of augmentation strategies on two personalization methods: DreamBooth and InstantID. We compare classical augmentations (flipping, cropping, color adjustments) with generative augmentation using InstantID’s synthetic images to enrich training data. Using SDXL and a new FaceDistance metric based on FaceNet, we quantitatively assess facial similarity. Results show classical augmentations can cause artifacts harming identity retention, while InstantID improves fidelity when balanced with real images to avoid overfitting. A user study with 97 participants confirms high photorealism and preferences for InstantID’s polished look versus DreamBooth’s identity accuracy. Our findings inform effective augmentation strategies for personalized text-to-image generation.
{strip}[Uncaptioned image]
Figure 1: Pipeline for generating personalized portraits using synthetic images enhanced through classical and generative augmentations to improve identity resemblance in DreamBooth and InstantID outputs.
@@ 1 Introduction
In this work, we analyze how augmentations affect personalized SD models trained via few-shot and zero-shot methods, focusing on DreamBooth and InstantID. We examine their impact on model performance and whether they improve the realism and consistency of generated images.
In particular, we analyze both classical and generative augmentation strategies to bridge the gap between limited real data and high-fidelity synthetic outputs. By refining facial features and preserving identity through targeted GenAI-based augmentations, such as InstantID, we aim to improve the applicability of personalized generation in scenarios where synthetic data must closely mirror real-world characteristics. We analyze under which conditions we can ensure that “GenAI outputs improve GenAI outputs”, avoiding a data quality collapse, providing best practices and heuristics.
@@ 2.2 Subject-Driven Image Generation
InstantID [24] is a zero-shot method combining facial feature extraction and text conditioning. It uses 5 facial landmarks to control the face’s position and orientation, providing greater control over the output.
We standardize experiments by using the same SDXL model for both methods to ensure a fair comparison.
@@ 2.3.1 Classical Image Augmentations
Classical image augmentations include geometric transformations (flipping, rotation, scaling, cropping), photometric adjustments (brightness, contrast, saturation, hue), and noise injections (Gaussian, salt-and-pepper). These predefined techniques preserve semantic integrity while adding controlled diversity to expand limited training data.
@@ 2.4 Face Processing Approaches
FaceNet [21] maps facial images to a 128-dimensional embedding space, where similar faces are positioned closely together. The standard pipeline uses MTCNN [27] for face detection before embedding generation, with cosine distance metrics for similarity assessment [22].
@@
(a) Real Images
Figure 2: Example Improvement: DreamBooth Results Comparing Classical vs. Generative Augmentations for Subject Dataset Vacation-Anna. The model is prompted with default prompt with batchsize 4. The results are not cherry-picked to resemble the downstream application use. Although (b) is visually more interesting, the method in (c) is more consistent across many subject datasets.
@@ 3 Methodology
3.1 Subject Datasets
Our internal dataset includes 3 to 15 images per participant, with 10 participants. To maintain naturalistic data collection, we instructed them: “Can you send me portrait/selfie-style photos of your face in different places? The more different places, the better.” By avoiding strict guidelines, we ensured the images reflect realistic user behavior. Consequently, our findings align well with real-world data distributions, improving transferability and applicability. The names used to refer to subjects in this dataset are pseudonyms and do not correspond to real individuals.
Our internal dataset exhibits diverse environmental conditions, facial orientations, and image qualities, ensuring variability that reflects real-world scenarios. The images include varied backgrounds, lighting, and subject behaviors, making the dataset both representative and robust. For example, some images feature cluttered or irregular backgrounds (e.g., Baker-Zoe, Bottle-Hugo), while others are from controlled settings (Biometric-Kora). Gaze variation is present: Doctor-Nina looks away from the camera, and 3D-Gary shows dynamic head movements from video. Differences in appearance and accessories also appear, such as Farmer-Lisa wearing a helmet and Staircase-Judy wearing makeup. Lighting varies from well-lit (Vacation-Anna) to suboptimal (2024-Kora), enhancing realism. These factors make our dataset valuable for evaluating model performance in unconstrained, real-world conditions.
@@ 3.2 Dataset Augmentations
Classical Augmentations  Standard techniques include: (i) Random Horizontal Flip with $p\in\{0,\frac{1}{2},1\}$, and (ii) Color Jitter varying brightness, saturation (±$(\pm 5,\pm 15)$ and hue (±$(\pm 5\degree,\pm 15\degree)$.
Background Augmentation  We use U2-Net [18] for subject isolation, testing both base and human segmentation models. Backgrounds include flat colors, Wikimedia Patterns [5], and Studio Backdrops (handpicked simple pictures from BG-20k [13], combined with our internal dataset).
@@
Generative Augmentation  Using InstantID, we generate new subject images with prompts from dolphin 2.2.1 - Mistral 7B [3] and varied facial landmarks. Based on one or more input images of a person, we run them through the InstantID pipeline with augmented landmarks and LLM generated prompts. The landmarks are extracted from the input images and then slightly perturbed to preserve resemblance while introducing diversity. (Omitting this step results in limited expression diversity, with generated images often replicating the same facial expression—an effect clearly visible in the rightmost examples of Figure 6.) This process yields a synthetic dataset of personalized subject images, which is subsequently used for downstream DreamBooth training to enhance model specificity and subject fidelity.
3.3 Hardware, Software, and Hyperparameters
@@
The default prompt is “A professional headshot of a subject wearing a suit in a well-lit studio, DSLR.” Empirically, including gender as man or woman imparts western cultural gender traits, which we prefer.
@@ 3.4 FaceDistance Metric
To quantify facial similarity in generated images, we employ the FaceDistance metric based on FaceNet embeddings [21]. FaceNet projects facial images into a 128-dimensional hyperspherical embedding space where spatial proximity reflects facial similarity.
Given a dataset of subject images, we first compute their FaceNet embeddings and derive the mean face vector, denoted as $\bar{v}_{\text{real}}$. For any generated image, we then calculate the cosine distance between its embedding vector and $\bar{v}_{\text{real}}$. A lower cosine distance implies greater resemblance to the real subject. FaceDistance serves as a practical tool for:
  • Coarsely ranking generated images by similarity (lower distances are better),
  • Discarding the top $k$of distant embeddings to improve personalization quality [^1] Empirical evidence supports $k=15$ for datasets with $n\geq 8$; independently, values above the 80th percentile $(q_{0.80})$ may be discarded.,
  • Identifying failure cases such as off-subject generations or severe artifacts.
In our experiments, the cosine distance between real images and $\bar{v}_{\text{real}}$ (within a person’s dataset) has a mean of $0.13$, a maximum of $0.22$, and a minimum of $0.05$.
For a detailed discussion of how FaceDistance enhances downstream applications such as large-scale image filtering and user satisfaction, see Suppl. 10.
@@ 4 Experiment Results
Despite using a realistic image generation model, achieving photorealistic face generation remains challenging without strict constraints on subject images. We relaxed many constraints to improve usability, since expecting average users to compile datasets without understanding image generation is difficult. High subject fidelity is crucial for effectiveness in downstream tasks, as humans are more sensitive to facial feature variations than to textures.
A major issue with unconstrained datasets is that the images may poorly represent the person. This is similar to struggling to recognize someone in real life based only on a few photographs. We observe this with small datasets where $\operatorname{size}\leq 3$. In such cases, generated images reflect the dataset well—outsiders unfamiliar with the person often consider the images accurate. However, those who know the person find the generated images to be poor representations.
@@ 4.1 DreamBooth
We configure our hyperparameters to prioritize high facial fidelity over recontextualization capabilities. Identity preservation is challenging in DreamBooth, so we prefer to overfit for strong subject fidelity, accepting limited generation diversity.
@@
The common theme in augmentations is that if the augmented image has any kind of artifact/anomaly, then the rare token will be associated with it. The supporting observations are (i) When background is replaced with a geometric pattern (from wikimedia patterns [26]), the model will focus on learning the pattern than the subject (ii) When image is upscaled with ESRGAN, the texture ESRGAN introduces say present in generations (iii) the masks generated with U2-Net is not pixel-perfect. and results in a mix around hair/air boundary. This mix becomes associated with the subject. The human segmentation models training data was not highly accurate around hairs but was better in identifying body parts. The base model is performs better around hair and was overall better. The robustness of human segmentation model is not needed. (iii) any kind of color jitter is visible in generated images. For example the saturation change of $0.1$ is clearly present in generations. (iv) using Adobe Lightroom as a preprocessing step resulted in better color graded generations compared to non-preprocessed datasets. (v) datasets with low contrast (e.g. exclusively Polaroid pictures) resulted in copying the photography style/lighting from the pictures — though this can be attributed to our hyperparameter configuration.
@@
Because of the low recontextualization capabilities, backgrounds becomes highly associated with the rare token. Replacing the background with Pastel Colors [               ] and Rainbow Colors [           ] led to eccentric and often unrealistic images, with the latter occasionally generating pictures without subjects. Gray (range [0–255]) offered the highest resemblance to the subject, while Dark Gray (range [0–127]) caused the model to disassociate the subject from its context. Because of problems with U2-Net, Light Gray (range [128–255]) background outperformed Dark Gray, especially in bright environments. We believe replacing all backgrounds with similar themed (colorful/gray) background degraded models recontextualisation capability. Wikimedia Patterns slowed down learning and degraded the image quality across all generations. Lastly, Studio Backdrops introduced irregularities that reduced the quality of the generated images which can be thought as similar to Wikimedia Patterns because backdrops has patterns. The regular patterns and backdrops captured models focus more than the subject, which is undesirable when the goal is to teach the subjects face.
@@
Random Horizontal Flip slowed learning due to face asymmetry, which confused facial features. Random Rotation introduced unnatural alignments and visual artifacts like objects or black padding bars. Color Jitter caused erratic generations, as brightness, contrast, saturation, and hue changes were linked to rare tokens. (Figure 3)
(a) Hue Shift
(b) Random Rotation
(c) Saturation Jitter
Figure 3: Generated Artifacts from Augmentations. Color jitter (a, c) introduces unnatural hues and saturation; rotation (b) distorts alignment and adds padding objects.
@@
Both Alpha Blending and Poison Blending are discouraged, as they require careful manual processing to achieve good results. These techniques are not straightforward to apply and can lead to undesirable artifacts if mishandled.
@@
Images around 1 megapixel (MP) performed best, providing a balanced resolution for high-quality generation. We believe this is due to SDXL’s training resolutions (see Appendix I of [17]) being $\approx$ 1 MP, and the model tends to perform poorly outside this range where it lacks sufficient training data. Upscaling with ESR-GAN introduced visible artifacts, especially around facial features. Upscaling with Lanczos was effective, particularly when starting from larger images. However, if the initial dataset contained low-resolution images, the generated images exhibited facial blurring due to the nature of the Lanczos algorithm. The difference between bicubic and Lanczos was negligible. Downscaling resulted in more pixelated generations compared to using original-sized images. It should be noted that our testing output resolution was $1024\times 1024$. Figure 4 demonstrates these findings.
(a) User Provided Dataset
(b) ESR-GAN Upscaled Dataset
(c) Downsampled Dataset
(d) Down then Upsampled Dataset
Figure 4: Comparison of generated image quality when conditioned on subjects from datasets with different resolution manipulations. The subfigure captions indicate the type of conditioning dataset used for generation, while the images show the resulting generated samples. (a) Original dataset images ($\approx$ 0.7MP). (b) Upsampling artifacts, notably around the eyes. (c) Simulated low-res input via 0.5$\times$ Lanczos downsampling; compare with (a). (d) Downsampled (0.5$\times$) then upsampled (2$\times$) with bicubic interpolation—results show smoothed textures.
@@
InstantID Augmentation Datasets augmented with InstantID yield clearly superior performance. The added images need to be diverse (i.e., generated with various text conditioning and different keypoint images). Since we trade recontextualization abilities for increased facial similarity, generating the same person in similar contexts is beneficial. DreamBooth achieves similar facial similarity compared to InstantID but allows for greater control. The rigidity caused by the keypoint images is eliminated. 2-step generation method is twice as computationally expensive than using InstantID once and is not practical as face replacement when augmenting for Dreambooth finetuning usage. In such scenarios, automated face replacement emerges as a more efficient and suitable alternative. Additionally, achieving proper prompt diversity can be challenging. We preferred InstantID over DreamBooth.
@@
The ratio of real to InstantID-generated images depends entirely on the diversity of the generated images. No single concept should comprise more than $25\%$ of the dataset. For example, if images labeled as “a [$V$] man in a library” exceed $25\%$, DreamBooth training will associate the rare token with the concept. This results in a final DreamBooth model that is unusable due to a complete loss of recontextualization ability caused by overfitting.
@@
Since InstantID generations are highly realistic, additional images can be generated to better represent the subject during DreamBooth training. Using the same diffusion model for both InstantID and DreamBooth helps integrate the subject effectively without changing its context, keeping the dataset distribution closer to the diffusion model’s generation space.
@@ 4.1.1 FaceDistance
We selected the “best” DreamBooth checkpoint by generating images of “a [$V$] man” in different contexts for all checkpoints and ranking them using FaceDistance. This discarded obviously bad checkpoints (e.g. anomalies in generations, unable to generate the subject, divergence) but is not able to rank “good enough” checkpoints within themselves. (Figure 5) The FaceNet manifold isn’t sensitive to very similar looking people. For a given hyperparameter configuration, a few tests show when the model will be converged to its best state (usually between 3k and 6k steps) and since FaceDistance isn’t able to differentiate between them, FaceDistance isn’t a useful tool for this purpose.
@@
Figure 5: Face Distance Distributions Across Finetuning Steps for SDXL Real Using Closeup-Kora as Subject. Kernel density estimates (KDEs) of face distances between 2,000 generated images and subject images are shown for different model checkpoints. The y-axis represents the image set generated by each model version, ordered by finetuning progress from “Base SDXL” (blue, prompt: “a man”) to finetuned checkpoints labeled “Ckpt N Steps” (red, prompt: “a [$V$] man”, where “[$V$]” denotes the rare token). The top row (green, “Subject Images”) shows self-similarity of the subject images. FaceDistance ($x$-axis) measures identity similarity, with lower values indicating closer matches. The mode of each distribution is marked with a yellow dot.